OpenStudy (anonymous):

Find w(-3) using synthetic substitution:; 11x^3 -6x^2 +2

OpenStudy (anonymous):

I don't know about synthetic substitution but I know other method: Put x = -3 here: \[11(-3)^3 - 6(-3)^2 + 2 \implies ??\]

OpenStudy (anonymous):

it would be: \[11 \times (-27) - 6 \times 9 + 2 = -349\]

OpenStudy (anonymous):

thanks